# From Continual Learning to SGD and Back: Better Rates for Continual Linear Models

### v2

@@ 6.2 A Universal $\mathcal{O}(1/\sqrt[4]{k})$ Rate for Random Orderings in Continual Linear Classification
Regularization methods are commonly used to prevent forgetting in CL [see 33]. Evron et al. [16] studied a regularized linear model for continual classification. They considered $T\geq 2$ jointly separable, binary classification tasks, defined by datasets $S_{1},\dots,S_{T}$ consisting of vectors ${\mathbf{x}\in\mathbb{R}^{d}}$ and their labels $y\in\left\{-1,+1\right\}$. They proved that a weakly-regularized scheme implicitly applies sequential max-margin projections. That is, in the limit as $\lambda\to 0$, the iterates of the two following schemes align in direction, enabling the study of continual classification through projection algorithms.
Scheme 5 Regularized Continual Classification
@@
Scheme 6 Sequential Max-Margin Projections
  Initialize $\mathbf{w}_{0}=\mathbf{0}_{d}$
   For each iteration $t=1,\dots,k$:
       $\mathbf{w}_{t}\leftarrow\bm{\Pi}_{\tau(t)}(\mathbf{w}_{t-1})\triangleq\operatorname*{argmin}_{\mathbf{w}\in\mathcal{C}_{\tau(t)}}\left\|{\mathbf{w}-\mathbf{w}_{t-1}}\right\|$ where $\mathcal{C}_{m}\!\triangleq\!\left\{\mathbf{w}\!\in\!\mathbb{R}^{d}\mid y\mathbf{w}^{\top}\mathbf{x}\geq 1,~{}\forall(\mathbf{x},y)\!\in\!S_{m}\right\}$
They studied forgetting under several orderings, using an equivalent of our Definition 2.4:
@@
The combination of our POCS rate (Theorem 6.1) and SGD stability arguments leads to the following.
@@
Under a random ordering, with or without replacement, over $T$ jointly separable tasks, the expected forgetting of the weakly-regularized Scheme 5 (at $\lambda\to 0$) after $k\geq 1$ iterations is bounded as,
$\mathbb{E}_{\tau}\big{[}F_{\tau}({k})\big{]}\leq\frac{7\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{\sqrt[4]{k}}\,,\quad\text{where }\mathbf{w}_{\star}\in{\operatorname*{argmin}}_{\mathbf{w}\in\mathcal{C}_{1}\cap\dots\cap\mathcal{C}_{T}}\left\|{\mathbf{w}_{0}-\mathbf{w}}\right\|^{2}\,.$
@@
Table 2: Forgetting Rates in Weakly-Regularized Continual Linear Classification on Separable Data. In all cells, we omit mild constant multiplicative factors and normalize by an unavoidable $\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}$ term.
@@
Evron et al. [16] $\exp\Big{(}{-\frac{k}{4T\|{\mathbf{w}_{\star}}\|^{2}R^{2}}}\Big{)}$ $\frac{T^{2}}{\sqrt{k}}\wedge\exp\Big{(}{-\frac{k}{16T^{2}\|{\mathbf{w}_{\star}}\|^{2}R^{2}}}\Big{)}$
Ours (2025) $\frac{1}{\sqrt[4]{k}}$ $\frac{1}{\sqrt[4]{T}}$
@@ 7 Discussion
Our work established a reduction from continual linear regression to the (block) Kaczmarz method and then to “stepwise-optimal” SGD. This enabled the development of analytic tools for last-iterate SGD schemes, leading to significantly improved and even universal rates for random orderings in continual learning and the Kaczmarz method. Our main results are summarized in Tables 1 and 3.
@@
Many popular continual benchmarks in deep learning implicitly assume a random ordering, such as the permuted MNIST benchmark [33]. Our paper shows that in sufficiently long task sequences, random ordering is enough to prevent catastrophic forgetting, and the training loss goes to zero, even in the worst case. In accordance with our results, Lesort et al. [39] examined a random CL benchmark—in which a subset of classes is randomly sampled in each task—and observed that forgetting diminishes as more tasks are sampled, even while training with standard SGD (without any modifications to mitigate forgetting). This suggests that random orderings may contaminate continual learning benchmarks, making it harder to isolate the algorithmic effects being tested. Furthermore, real-world tasks often change gradually, not adhering to random orderings. Such “gradually evolving” datasets might be more challenging and relevant as continual benchmarks.
@@
In Section 3.1 we revisited known connections between continual regression and the Kaczmarz method [15]. We broadened this connection in Section 3.2, bridging the block Kaczmarz method and “stepwise-optimal” SGD, thus applying our novel SGD bounds to the Kaczmarz method. Using Kaczmarz terminology, given a system $\mathbf{A}\mathbf{x}=\mathbf{b}$ consisting of $T$ blocks of an average rank $\bar{r}$ where $\mathbf{A}_{m}\in\mathbb{R}^{{n_{m}}\times d},\mathbf{b}_{m}\in\mathbb{R}^{n_{m}}$, our rates from Section 4 can be summarized as $\mathbb{E}_{\tau}\big{[}\frac{1}{2T}\sum_{m=1}^{T}\big{\|}{\mathbf{A}_{m}\mathbf{x}_{k}-\mathbf{b}_{m}}\big{\|}^{2}\big{]}=\mathcal{O}\big{(}{\min\big{(}{k^{-1/4},\frac{1}{k}\sqrt{d-\bar{r}},\frac{1}{k}\sqrt{T\bar{r}}}\big{)}}\big{)}$ for random orderings with replacement and $\mathcal{O}\big{(}{\min\big{(}{k^{-1/4},\frac{1}{k}\left({d-\bar{r}}\right)}\big{)}}\big{)}$ without replacement. Note that we bounded the loss, rather than the “error” $\left\|{\mathbf{w}_{k}-\mathbf{w}_{\star}}\right\|^{2}$, thus enabling the derivation of rates independent of quantities like the condition number that can make convergence arbitrarily slow.
@@
The seminal work of Strohmer and Vershynin [66] proposed a Kaczmarz variant that samples rows with probability proportional to their squared norm. Our approach accommodates non-uniform sampling, including norm based ones. This may tighten Theorem 4.2, replacing the dependence on the maximum row norm $R$ with the average one. In the block version, both uniform and non-uniform variants exist [50, 21]. There also, our approach should yield better bounds when weighting blocks by to their norms, alleviating the dependence on the maximum norm.
Future Work.
We narrowed the gap between existing lower and upper worst-case bounds for random orderings in continual linear regression (see Table 1). However, a considerable gap remains between $\Omega(1/k)$ and $\mathcal{O}({1/k^{1/4}})$. Generally, we conjecture that the last-iterate SGD rates can be improved beyond those in Theorem 5.1, and that Theorem E.4 can be extended to the multi-epoch setup. Following our reductions (Section 3), improved rates for “stepwise-optimal” SGD rates would immediately refine the bounds for continual linear regression and classification.
Acknowledgments and Disclosure of Funding
@@ Appendix A Related Work
For the general (non-realizable) smooth stochastic setup, the recent work of Liu and Zhou [43] was the first (and only, to our knowledge) to provide upper bounds on the convergence rate of the last SGD iterate. While their bounds are applicable in the realizable setting, they require non-constant step sizes to obtain non-trivial convergence, and are therefore not useful for our purposes (see Table 3). Our analysis technique in Section 5 borrows from the work of Shamir and Zhang [62, also mentioned in Table 3] which, in fact, belongs to the comparatively-richer line of work on the non-smooth setting [62, 30, 71, 43]. Notably, SGD in a stochastic non-realizable (either smooth or non-smooth) setup requires uniformly bounded noise assumptions, and generally cannot accommodate a constant step size independent of the optimization horizon.
@@
Stochastic (*) Shamir and Zhang [62] $\frac{1}{\eta T}+\eta G^{2}\log T$
Deterministic Smooth ($\sigma=0$) Nesterov [52] $\frac{1}{(2-\eta\beta)\eta T}$
Stochastic Smooth Lan [37] $\frac{1}{\eta T}+\eta\sigma^{2}$
Liu and Zhou [43] $\frac{1}{\eta T}+\eta\sigma^{2}\log T$
Stochastic Smooth Realizable ($\sigma_{\star}=0$) Srebro et al. [65] $\frac{1}{(1-\eta\beta)\eta T}$
Stochastic Regression Realizable ($\sigma_{\star}=0$) Bach and Moulines [1] $\frac{1}{\eta T}$
Varre et al. [67] $\frac{1}{(1-2\eta\beta\log T)\eta T}$
Ours (2025)  $\frac{1}{(2-\eta\beta)\eta T^{1-\eta\beta(1-\eta\beta/4)}}$
(*) They consider bounded domains [62]; Orabona [53], Liu and Zhou [43] obtain similar bounds for the unconstrained case. For non-fixed step sizes, Jain et al. [30] obtain minimax optimal bounds without log factors.
Our analysis for SGD without-replacement is related to a long line of work primarily focused on the average iterate convergence rates [e.g., 58, 48, 59, 56, 47, 11, 10]. For the non-strongly convex case, near-optimal bounds (for the average iterate) have been established for the general smooth case [48, 47]. In a subsequent work, Cai et al. [10] refined the dependence on problem parameters for the smooth realizable case (among others). Guarantees for the last iterate have only been established recently by Liu and Zhou [44] and Cai and Diakonikolas [9]. However, their bounds decay with the number of epochs rather than the number of iterations and apply only to non-constant step sizes, making them inapplicable to our setting. Specifically, in a realizable $\beta$-smooth setup, after $J$ without-replacement SGD epochs over a finite sum of size $n$, Mishchenko et al. [47], Cai et al. [10] obtained an $O(\beta/J)$ bound for the average iterate with step size $\eta=1/(\beta n)$; and Liu and Zhou [44], Cai and Diakonikolas [9] derived similar bounds for the last iterate up to logarithmic factors.
Another line of work related to ours studies algorithmic stability [6, 61] of gradient methods, which is the main technique we use in the proof of Theorem E.4. Our approach is similar in nature to that of Nagaraj et al. [48], Sherman et al. [63], Koren et al. [35] and primarily builds on Sherman et al. [63], who were the first to formally introduce the notion of without-replacement stability. For with-replacement SGD, [25] discussed its algorithmic stability under smooth loss functions. Later, [38], improved this bound in the realizable loss case. The case we consider—i.e., the stability of without-replacement SGD under smooth and realizable loss functions—is not covered in the existing literature.
@@
Our results in Section 4 establish universal bounds for random orderings, both with and without replacement. Both the with- and without-replacement variants converge linearly towards the minimum-norm solution $\mathbf{w}_{\star}$ [21, 24], but as we explained in Section 7, the rates can be arbitrarily slow. Recht and Ré [57] formulated a noncommutative analog of the arithmetic-geometric mean inequality that, if true, could have shown that without-replacement orderings lead to faster loss convergence than with-replacement orderings in Kaczmarz methods, and consequently in continual linear regression. Years later, Lai and Lim [36] proved that this inequality does not hold in general [see also 12]. Moreover, as in other areas, empirical studies found that row shuffling followed by cyclic orderings performs as well as i.i.d. orderings [54]. This naturally connects to interesting observations and open questions regarding various forms of shuffled SGD [5, 70]. Our rates are similar for both with- and without-replacement orderings (up to small constants), meaning they do not indicate a clear advantage for either. However, we believe they are far from tight, leaving interesting open questions in this direction.
@@
The NLMS algorithm is a classical adaptive filtering method. In its simplest version [64], the method perfectly fits a single—usually noisy—random sample at a time, using the same update rule as the Kaczmarz method (and thus, as our continual Scheme 1 in a rank-$1$ case). There also exists a more complex version of this method, which uses more samples per update [60]. Both papers give strong $\mathcal{O}(1/k)$ MSE rates in the noiseless setting (matching our realizable setting). However, they assume a very limited data model, where the sampled vectors are either orthogonal or identical up-to-scaling. Under such conditions, Evron et al. [15] showed that there is no forgetting (of previously learned tasks), implying that the MSE decays as the number of tasks still unseen at time $k$.

### v3

@@ 6.2 A Universal $\mathcal{O}(1/\sqrt[4]{k})$ Rate for Random Orderings in Continual Linear Classification
Regularization methods are commonly used to prevent forgetting in CL (see Kirkpatrick et al., 2017; Aljundi et al., 2018; Li et al., 2023). Evron et al. (2023) studied a weakly-regularized linear model for continual classification. They considered $T\geq 2$ jointly separable, binary classification tasks, defined by datasets $S_{1},\dots,S_{T}$ consisting of vectors $\mathbf{x}\in\mathbb{R}^{d}$} and their labels $y\in\left\{-1,+1\right\}$.
Scheme 5 Weakly-Regularized Continual Linear Classification (for $\lambda\to 0$)
@@
Specifically, Evron et al. (2023) proved that learning an entire (separable) classification task in this continual scheme implicitly applies projection onto convex sets. More formally,
Reduction 4 (Continual Classification $\Rightarrow$ POCS)
Given jointly separable tasks, the continual iterates of Scheme 5 in the limit as $\lambda\to 0$, align in direction of the sequential projections of Scheme 4 onto the convex sets defined as $\mathcal{C}_{m}\triangleq\left\{\mathbf{w}\in\mathbb{R}^{d}\mid y\mathbf{w}^{\top}\mathbf{x}\geq 1,~\forall(\mathbf{x},y)\in S_{m}\right\},\,\forall m\in\left[{T}\right]$.
Importantly, this reduction enables the study of continual classification through projection algorithms. In particular, Evron et al. (2023) studied forgetting using an equivalent of our Section 2:
@@
As illustrated in Figure 2, we derive the following bound by combining their reduction with our POCS rate (Theorem 6.1) and SGD stability arguments.
@@
Under a random ordering, with or without replacement, over $T$ jointly separable tasks, the expected forgetting of the weakly-regularized Scheme 5 after $k\geq 1$ iterations is,
$\mathbb{E}_{\tau}\!\left[F_{\tau}({k})\right]\leq\frac{7\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{\sqrt[4]{k}}\,,\quad\text{where }\mathbf{w}_{\star}\in{\operatorname*{argmin}}_{\mathbf{w}\in\mathcal{C}_{1}\cap\dots\cap\mathcal{C}_{T}}\left\|{\mathbf{w}_{0}-\mathbf{w}}\right\|^{2}\,.$
@@
Table 2: Forgetting Rates in Weakly-Regularized Continual Linear Classification on Separable Data. All cells omit mild multiplicative constants and normalize by an unavoidable $\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}$ term.
@@
  Evron et al. (2023) $\exp\Big({-\frac{k}{4T\|{\mathbf{w}_{\star}}\|^{2}R^{2}}}\Big)$ $\frac{T^{2}}{\sqrt{k}}\wedge\exp\Big({-\frac{k}{16T^{2}\|{\mathbf{w}_{\star}}\|^{2}R^{2}}}\Big)$
Ours $\frac{1}{\sqrt[4]{k}}$ $\frac{1}{\sqrt[4]{T}}$
@@ 7 Discussion
Our work established reductions from continual linear regression and classification to “stepwise-optimal” SGD. This enabled the development of analytic tools for last-iterate SGD schemes, leading to significantly improved and even universal rates for random orderings in continual learning. Our main results are summarized in Tables 1, 2 and 3.
@@
Many popular continual benchmarks in deep learning implicitly assume a random ordering, such as the permuted MNIST benchmark (Kirkpatrick et al., 2017). Our paper shows that in sufficiently long task sequences, random ordering is enough to prevent catastrophic forgetting, and the training loss goes to zero, even in the worst case. In accordance with our results, Lesort et al. (2023) examined a random CL benchmark—in which a subset of classes is randomly sampled in each task—and observed that forgetting diminishes as more tasks are sampled, even while training with standard SGD (without any modifications to mitigate forgetting). This suggests that random orderings may contaminate continual learning benchmarks, making it harder to isolate the algorithmic effects being tested. Furthermore, real-world tasks often change gradually, not adhering to random orderings. Such “gradually evolving” datasets might be more challenging and relevant as continual benchmarks.
@@
In Section 3.1 we revisited known connections between continual regression and the Kaczmarz method (Evron et al., 2022). We broadened this connection in Section 3.2, bridging the block Kaczmarz method and “stepwise-optimal” SGD, thus applying our novel SGD bounds to the Kaczmarz method. Using Kaczmarz terminology, given a system $\mathbf{A}\mathbf{x}=\mathbf{b}$ consisting of $T$ blocks of an average rank $\bar{r}$ where $\mathbf{A}_{m}\in\mathbb{R}^{{n_{m}}\times d},\mathbf{b}_{m}\in\mathbb{R}^{n_{m}}$, our rates from Section 4 can be summarized as $\mathbb{E}_{\tau}\big[\frac{1}{2T}\sum_{m=1}^{T}\big\|{\mathbf{A}_{m}\mathbf{x}_{k}-\mathbf{b}_{m}}\big\|^{2}\big]=\mathcal{O}\big({\min\big({k^{-1/4},\frac{1}{k}\sqrt{d-\bar{r}},\frac{1}{k}\sqrt{T\bar{r}}}\big)}\big)$ for random orderings with replacement and $\mathcal{O}\big({\min\big({k^{-1/4},\frac{1}{k}\left({d-\bar{r}}\right)}\big)}\big)$ without replacement. Note that we bounded the loss, rather than the “error” $\left\|{\mathbf{w}_{k}-\mathbf{w}_{\star}}\right\|^{2}$, thus enabling the derivation of rates independent of quantities like the condition number that can make convergence arbitrarily slow.
@@
The seminal work of Strohmer and Vershynin (2009) proposed a Kaczmarz variant that samples rows with probability proportional to their squared norm. Our approach also accommodates non-uniform sampling, including norm-based ones. Specifically, Claim 1 tightens Theorem 4.3 by employing norm-weighted block sampling, thereby replacing the dependence on the maximum row norm $R$ with the average $\bar{R}$.
From Training to Generalization Results.
The rates derived in our paper apply to the forgetting and loss on the training sets (Sections 2 and 2.2). These rates can be extended to the generalization loss at the cost of an additive $\mathcal{O}(1/\sqrt{N})$ term (where $N$ is the number of samples per task) via uniform-convergence arguments—or even an additive $\mathcal{O}(1/N)$ term using the more refined Rademacher bounds for linear models from Srebro et al. (2010a).
Future and Follow-up Work.
We narrowed the gap between existing lower and upper worst-case bounds for random orderings in continual linear regression (see Table 1). However, a considerable gap remains between $\Omega(1/k)$ and $\mathcal{O}({1/k^{1/4}})$, largely stemming from our proof technique in Lemma 5.3. Our argument follows the approach of Shamir and Zhang (2013), originally developed to control the last SGD iterate in the convex, non-smooth regime. When instantiated in our (smooth, realizable) setting, this approach introduces an exponential dependence on $\eta\beta$. Because of this exponential sensitivity, employing coarse inequalities (e.g., bounding a negative term by 0) can be costly: even small constant-factor losses may effectively change the power of $T$ appearing in the final bound.
Following our reductions (Sections 3 and 6), improved rates for “stepwise-optimal” SGD rates would immediately refine the bounds for continual linear regression and classification. Indeed, a follow-up work by Attia et al. (2025) departs from Shamir and Zhang (2013) and builds on a different technique proposed by Zamani and Glineur (2023). Using a more refined analysis, they establish an improved rate for SGD and, combined with our reductions, obtain a tighter upper bound of $\mathcal{O}(1/\sqrt{k})$ for continual linear regression (and classification). Finding the exact worst-case complexity between $\Omega(1/k)$ and $\mathcal{O}({1/\sqrt{k}})$ remains an open question.
Finally, we note that our analytical flow relies on Reduction 1, which in turn assumes that the continual Scheme 1 learns each task to convergence. Nonetheless, a concurrent follow-up by Levinstein et al. (2025) follows a similar overall analytical flow, despite studying more practical continual learning schemes that do not learn to convergence—namely, those using $\ell_{2}$ regularization or finite step budgets. Similarly to us, they reduce learning an entire task to a single gradient step and apply last-iterate SGD analysis to obtain convergence rates.
[…]
@@ Appendix A Related Work
For the general (non-realizable) smooth stochastic setup, the recent work of Liu and Zhou (2024a) was the first (and only, to our knowledge) to provide upper bounds on the convergence rate of the last SGD iterate. While their bounds are applicable in the realizable setting, they require non-constant step sizes to obtain non-trivial convergence, and are therefore not useful for our purposes (see Table 3). Our analysis technique in Section 5 borrows from the work of Shamir and Zhang (2013, also mentioned in Table 3) which, in fact, belongs to the comparatively-richer line of work on the non-smooth setting (Shamir and Zhang, 2013; Jain et al., 2019; Zamani and Glineur, 2023; Liu and Zhou, 2024a). Notably, SGD in a stochastic non-realizable (either smooth or non-smooth) setup requires uniformly bounded noise assumptions, and generally cannot accommodate a constant step size independent of the optimization horizon.
@@
  Stochastic (*) Shamir and Zhang (2013) $\frac{1}{\eta T}+\eta G^{2}\log T$
Deterministic Smooth ($\sigma=0$) Nesterov (1998) $\frac{1}{(2-\eta\beta)\eta T}$
Stochastic Smooth Lan (2012) $\frac{1}{\eta T}+\eta\sigma^{2}$
Liu and Zhou (2024a) $\frac{1}{\eta T}+\eta\sigma^{2}\log T$
Stochastic Smooth Realizable ($\sigma_{\star}=0$) Srebro et al. (2010b) $\frac{1}{(1-\eta\beta)\eta T}$
Stochastic Regression Realizable ($\sigma_{\star}=0$) Bach and Moulines (2013) $\frac{1}{\eta T}$
Varre et al. (2021) $\frac{1}{(1-2\eta\beta\log T)\eta T}$
Ours $\frac{1}{(2-\eta\beta)\eta T^{1-\eta\beta(1-\eta\beta/4)}}$
(*) They consider bounded domains (Shamir and Zhang, 2013); Orabona (2020); Liu and Zhou (2024a) obtain similar bounds for the unconstrained case. For non-fixed step sizes, Jain et al. (2019) obtain minimax optimal bounds without log factors.
Our analysis for SGD without-replacement is related to a long line of work primarily focused on the average iterate convergence rates (e.g., Recht and Ré, 2012b; Nagaraj et al., 2019; Safran and Shamir, 2020; Rajput et al., 2020; Mishchenko et al., 2020; Cha et al., 2023; Cai et al., 2024). For the non-strongly convex case, near-optimal bounds (for the average iterate) have been established for the general smooth case (Nagaraj et al., 2019; Mishchenko et al., 2020). In a subsequent work, Cai et al. (2024) refined the dependence on problem parameters for the smooth realizable case (among others). Guarantees for the last iterate have only been established recently by Liu and Zhou (2024b) and Cai and Diakonikolas (2025). However, their bounds decay with the number of epochs rather than the number of iterations and apply only to non-constant step sizes, making them inapplicable to our setting. Specifically, in a realizable $\beta$-smooth setup, after $J$ without-replacement SGD epochs over a finite sum of size $n$, Mishchenko et al. (2020); Cai et al. (2024) obtained an $O(\beta/J)$ bound for the average iterate with step size $\eta=1/(\beta n)$; and Liu and Zhou (2024b); Cai and Diakonikolas (2025) derived similar bounds for the last iterate up to logarithmic factors.
Another line of work related to ours studies algorithmic stability (Bousquet and Elisseeff, 2002; Shalev-Shwartz et al., 2010) of gradient methods, which is the main technique we use in the proof of Theorem E.9. Our approach is similar in nature to that of Nagaraj et al. (2019); Sherman et al. (2021); Koren et al. (2022) and primarily builds on Sherman et al. (2021), who were the first to formally introduce the notion of without-replacement stability. For with-replacement SGD, Hardt et al. (2016) discussed its algorithmic stability under smooth loss functions. Later, Lei and Ying (2020), improved this bound in the realizable loss case. The case we consider—i.e., the stability of without-replacement SGD under smooth and realizable loss functions—is not covered in the existing literature.
@@
Our results in Section 4 establish universal bounds for random orderings, both with and without replacement. Both the with- and without-replacement variants converge linearly towards the minimum-norm solution $\mathbf{w}_{\star}$ (Gower and Richtárik, 2015; Han and Xie, 2024), but as we explained in Section 7, the rates can be arbitrarily slow. Recht and Ré (2012a) formulated a noncommutative analog of the arithmetic-geometric mean inequality that, if true, could have shown that without-replacement orderings lead to faster loss convergence than with-replacement orderings in Kaczmarz methods, and consequently in continual linear regression. Years later, Lai and Lim (2020) proved that this inequality does not hold in general (see also De Sa, 2020). Moreover, as in other areas, empirical studies found that row shuffling followed by cyclic orderings performs as well as i.i.d. orderings (Oswald and Zhou, 2015). This naturally connects to interesting observations and open questions regarding various forms of shuffled SGD (Bottou, 2009; Yun et al., 2021). Our rates are similar for both with- and without-replacement orderings (up to small constants), meaning they do not indicate a clear advantage for either. However, we believe they are far from tight, leaving interesting open questions in this direction.
@@
The NLMS algorithm is a classical adaptive filtering method. In its simplest version (Slock, 1993), the method perfectly fits a single—usually noisy—random sample at a time, using the same update rule as the Kaczmarz method (and thus, as our continual Scheme 1 in a rank-$1$ case). There also exists a more complex version of this method, which uses more samples per update (Sankaran and Beex, 2000). Both papers give strong $\mathcal{O}(1/k)$ MSE rates in the noiseless setting (matching our realizable setting). However, they assume a very limited data model, where the sampled vectors are either orthogonal or identical up-to-scaling. Under such conditions, Evron et al. (2022) showed that there is no forgetting (of previously learned tasks), implying that the MSE decays as the number of tasks still unseen at time $k$.
Alternative Continual Schemes That Do Not Forget.
Schemes such as Recursive Least Squares (RLS; see Chong and Zak 2004, Chapter 12.2) and its block variant BRMP (Zhuang et al., 2021) provide analytical alternatives to the block Kaczmarz method (Scheme 2). These methods effectively avoid forgetting by maintaining an $\mathcal{O}(d^{2})$ matrix. See also Proposition 5.5 in Evron et al. (2023).
Importantly, we study a continual scheme (Scheme 1) that closely characterizes common training practices. In particular, training with (S)GD to convergence coincides with the analytical updates of the (block) Kaczmarz method (Scheme 2), making the latter illustrative of most gradient-based continual learning approaches. Understanding continual gradient-based algorithms in linear models is especially relevant given the linear dynamics of deep neural networks in the NTK regime (Jacot et al., 2018). These regimes are typically highly overparameterized, yet they remain covered by our analysis and by the dimensionality-independent rates we derive for the naturally forgetful but memoryless Scheme 1.